In a game, you have a 1/34 probability of winning $91 and a 33/34 probablility of losing $9. What is your expected value?
Answer by fractalier(2101)

(1/34)(91) + (33/34)(-9) = 91/34 - 297/34 = -206/34 = -103/17 = -6 1/17 dollars lost